HIGH-RESOLUTION PFI-ZEKE PHOTOELECTRON SPECTROSCOPY OF THE CYCLOPENTADIENYL RADICAL

\underline{H. J. WÖRNER} and F. MERKT, PHYSIKALISCHE CHEMIE, ETH ZÜRICH, CH-8093 ZÜRICH.

This presentation summarizes the results of the first photoelectron spectroscopic study of the cyclopentadienyl radical. The cyclopentadienyl radical was produced in a supersonic expansion by photolysis of cyclopentadiene at 248 nm using the method described in Ref. a. PFI-ZEKE photoelectron spectra were recorded using two different excitation schemes. In the first experiment single-photon excitation from the X̃ ^{2}E′′_{1} ground state of the radical was used to study the ground state of the cyclopentadienyl cation. In the second experiment, a resonance-enhanced two-photon excitation scheme via the vibrationless transition to the Ã ^{2}A′′_{2} studied in Ref. b provided access to the first electronically excited state of the cyclopentadienyl cation. The analysis of the spectra reveals that the ground state of the cyclopentadienyl cation is a triplet state of electronic symmetry ^{3}A′_{2} and the first excited state is a singlet of symmetry ^{1}E′_{2}, which is subject to a pronounced Jahn-Teller distortion. The results of ab initio calculations together with a preliminary analysis of the Jahn-Teller effect in the ^{1}E′_{2} state are presented.

a. S. Willitsch, J.M. Dyke and F. Merkt, Helv. Chim. Acta, 86, 1152 (2003). b. L. Yu, S.C. Foster, J.M. Williamson, M.C. Heaven and T.A. Miller, J. Chem. Phys. 92, 4263 (1988).